OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form. y = 3x – 6 y = 3x – 3 y = 3x + 3 y = 3x + 6

OpenStudy (anonymous):

gradient = (6-(-3))/(3-0) = 9/3 = 3 so it has to be 3x. Intercept happens at x=0 and they tell you one point is (0,-3) which is when x=0 therefore y= -3 which is intercept. So equation is y=3x -3

OpenStudy (anonymous):

y=3x-3

OpenStudy (anonymous):

y=mx+b where m is slop m=(y2-y1)/(x2-x1) m=3 y=3x+b (3,6) 6=3(3)+b b=-3 then y=3x-3

OpenStudy (anonymous):

Thank you <3

OpenStudy (anonymous):

u're welcome